OpenStudy (anonymous):

(-4+√(-16-4*1*3))/(2*1) I need this done ASAP!!! Aaagghh

OpenStudy (jdoe0001):

hmm use PEMDAS, what would you get in the radical?

OpenStudy (anonymous):

Me and my family have no clue what to do! We tried PEMDAS but we cant figure it out and its for school so i need to get it in today :P

OpenStudy (anonymous):

If you know, maybe you could walk me through it ?

OpenStudy (jdoe0001):

grouped elements first now inside them multiplication first FROM LEFT TO RIGHT sum/subtraction next, FROM LEFT TO RIGHT \(\bf \cfrac{-4+\sqrt{-16-\color{red}{4\times 1\times 3}}}{2\times 1}\)

OpenStudy (jdoe0001):

so... what does the multiplication gave you?

OpenStudy (anonymous):

in the red or at the bottom? Im so sorry im so math dumb. Ugh.

OpenStudy (jdoe0001):

the red part for now....lemme include the bottom \(\bf \bf \cfrac{-4+\sqrt{-16-\color{red}{4\times 1\times 3}}}{ \color{red}{2\times 1}} \)

OpenStudy (anonymous):

okay so the top red is 12 and the bottom is 2

OpenStudy (jdoe0001):

next the radical amounts then \(\bf \cfrac{-4+\sqrt{-16-4\times 1\times 3}}{2\times 1}\implies\cfrac{-4+\sqrt{\color{red}{-16-12}}}{2}\)

OpenStudy (anonymous):

4 and 3.4??

OpenStudy (jdoe0001):

well,.... -16 -12 is ... no ... 3.4 :|

OpenStudy (anonymous):

omg i thought something else lol sorry i mean -28

OpenStudy (anonymous):

Maybe..

OpenStudy (jdoe0001):